### v1

# Counterfactuals As a Means for Evaluating Faithfulness of Attribution Methods in Autoregressive Language Models

Abstract
Despite the widespread adoption of autoregressive language models, explainability evaluation research has predominantly focused on span infilling and masked language models (MLMs). Evaluating the faithfulness of an explanation method—how accurately the method explains the inner workings and decision-making of the model—is very challenging because it is very hard to separate the model from its explanation. Most faithfulness evaluation techniques corrupt or remove some input tokens considered important according to a particular attribution (feature importance) method and observe the change in the model’s output. This approach creates out-of-distribution inputs for causal language models (CLMs) due to their training objective of next token prediction. In this study, we propose a technique that leverages counterfactual generation to evaluate the faithfulness of attribution methods for autoregressive language modeling scenarios. Our technique creates fluent and in-distribution counterfactuals that makes evaluation protocol more reliable.
Counterfactuals As a Means for Evaluating Faithfulness of Attribution Methods in Autoregressive Language Models
Sepehr Kamahi1, Yadollah Yaghoobzadeh1,2 1School of Electrical and Computer Engineering College of Engineering, University of Tehran, Tehran, Iran 2Tehran Institute for Advanced Studies, Khatam University, Tehran, Iran sepehr.kamahi@ut.ac.ir, y.yaghoobzadeh@ut.ac.ir
1 Introduction
Most state-of-the-art NLP systems use autoregressive transformer-based language models (Touvron et al., 2023; Brown et al., 2020; Groeneveld et al., 2024). Since these models are opaque, there is great motivation to understand their decision-making process, and so, explanation methods are becoming increasingly important.
Attribution or feature importance (FI) is a common approach for explaining models (Zhao et al., 2023). Attribution methods’ objective is to explain which input features are most salient in the model’s predictions. These methods achieve this objective by assigning a scalar to each input feature, representing its significance to the model’s decision-making process. Input features in most NLP tasks are sub-words or some combination of them.
A pressing issue for testing and evaluating faithfulness of attribution methods is that most techniques focus on denoising language models or masked language models (Kobayashi et al., 2020, 2021; Ferrando et al., 2022b; Modarressi et al., 2022, 2023; Mohebbi et al., 2023), and most recent work dealing with autoregressive language models has focused on the plausibility of attributions (Yin and Neubig, 2022; Ferrando et al., 2023). Plausibility (or persuasiveness) might be an objective of the explainer, but the explainee’s objective is to understand the model’s decision-making process, not being persuaded that the decision is correct (Jacovi and Goldberg, 2021).
Almost all previous faithfulness evaluation techniques corrupt the input in one way or another, i.e., masking or erasing important tokens according to the attribution technique, and then looking at the change in the prediction. These techniques probably work fine for MLMs, which are pre-trained for mask and span infilling. However, for autoregressive language models like GPT-2, pre-trained for next token prediction, masking or erasing creates an out-of-distribution (OOD) input for the model. In this case, it is unclear whether corruption techniques evaluate the informativeness of the corrupted tokens or the robustness of the model to unnatural text and the artifacts introduced during test time (Hooker et al., 2019). Furthermore, inputs being OOD makes the explanation socially misaligned Hase et al. (2021); meaning what people expect from FI explanations—demonstrating which features are the most important in the model’s decision-making—and what the FI shows become different because FI becomes dependent on model priors rather than removal of features that the model has learned to rely on.
In this work, inspired by counterfactual generation—changing the model’s input in a way that flips the output—we develop a technique for evaluating the faithfulness of attribution methods in the autoregressive generation scenario. We use counterfactual generators to change the input, focusing on the important tokens specified by the attribution methods, and make sure that the input to the model is natural, fluent, and in-distribution. This way, we know that the change in the model’s prediction is because of changing the important tokens and not because of the input being out of distribution.
We argue that if an attribution method helps the counterfactual generator to change the model’s prediction with fewer changes, that method knows more about the model’s inner workings, which means it is more faithful. Also, because of the large output space of autoregressive language models like GPT-2 and LLaMA, including often thousands of vocabulary items, looking at the entire output space does not provide much insight. We use contrastive explanation Yin and Neubig (2022), which means looking only at the token predicted by the model and a foil token.
We apply our faithfulness evaluation approach to several attribution techniques including gradient norm, gradient $\times$ input, Erasure, KernelSHAP, and integrated gradient in the next word prediction of two LMs: our fine-tuned Gemma-2b and off-the-shelf Gemma-2b-instruct (Team et al., 2024).
Our contributions are as follows: (i) We introduce a novel faithfulness evaluation protocol, focused on maintaining the input data distribution of the model, suitable for the attribution methods of language models. (ii) We evaluate and rank several popular attribution methods using our approach. (iii) We demonstrate the difference in sensitivity between fine-tuned and off-the-shelf models when dealing with out-of-distribution (OOD) data and provide a solution.
2 Related work
Evaluating Explanations. Current faithfulness metrics mostly use removing important tokens or retraining only on important tokens identified by the attribution methods, (Chan et al., 2022). Abnar and Zuidema (2020) use agreement with gradient and ablation methods as an evaluation of their explanation methods. Wiegreffe and Pinter (2019) acknowledges that gradient methods should not be treated as ideal or ground truth nonetheless use the gradient as a proxy of the model’s intrinsic semantics. Explanations’ trustworthiness is task- and model-specific Bastings et al. (2022), and different attribution methods give deeply inconsistent results Neely et al. (2022). hence using one explanation method as the standard or the ground truth in all scenarios does not seem to be justifiable. DeYoung et al. (2020) introduce comprehensiveness, if only the chosen important tokens are used to make the prediction (are highlighted inputs necessary), and sufficiency, if the chosen important tokens on their own are sufficient to make the prediction. Carton et al. (2020) introduce a normalized version of comprehensiveness and sufficiency by dividing these measures into null difference. The Null difference is the sufficiency of an empty input or comprehensiveness of a full input. It is unclear whether corruption techniques evaluate the informativeness of the corrupted tokens or the robustness of the model to unnatural text and the artifacts introduced during test time.
Han et al. (2020) and Jain et al. (2020) consider an attribution method to be either faithful or unfaithful, with no middle ground, and use the term ’faithful by construction.’ In contrast, other works argue that faithfulness exists on a spectrum and that we should evaluate the ’degree of faithfulness’ of an explanation method (Jacovi and Goldberg, 2020). We adopt the latter approach, searching for a sufficiently faithful explanation method for our tasks.
Atanasova et al. (2023) use counterfactuals to evaluate faithfulness of natural language explanations—When we want the model to tell us in natural language why it made a particular decision—they use counterfactual generation techniques from (Ross et al., 2021).
Another line of work tries to evaluate explanations using uncertainty estimation. Slack et al. (2021) develop a Bayesian framework for generating feature importance estimates along with their associated uncertainty in the form of credible intervals.
OOD Problem in Explainability. Hooker et al. (2019) and Vafa et al. (2021) suggest retraining or fine-tuning the model with inputs that are partly erased to achieve a model that has the same train and evaluation distribution. However, this retraining is expensive and not always possible. Kim et al. (2020) try to solve the OOD problem by making the explanation itself in-distribution.
Feature Importance (Attribution). attributions are local explanations that assign a score to each input feature (token embeddings in most NLP tasks). The score represents how important that feature is for the predictor model according to that explanation method. We can categorize these methods into four types. i) Perturbation-based methods which work by perturbing the input examples such as removing or masking to measure feature importance (Li et al., 2016, 2017; Feng et al., 2018; Wu et al., 2020). ii) Gradient-Based methods determine the importance of each input feature by measuring the derivative of output with respect to each input (Mohebbi et al., 2021; Kindermans et al., 2019; Sundararajan et al., 2017; Lundstrom et al., 2022; Enguehard, 2023; Sanyal and Ren, 2021; Sikdar et al., 2021). iii) Surrogate-based methods use a simple, interpretable model to explain the original complex, black-box model (Ribeiro et al., 2016; Lundberg and Lee, 2017; Kokalj et al., 2021). iv) Decomposition-based methods try to break down the importance score into linear contributions from the input (Montavon et al., 2019; Voita et al., 2021; Chefer et al., 2021; Modarressi et al., 2022; Ferrando et al., 2022a).
3 Our method
Our faithfulness evaluation protocol consists of two models: the counterfactual generator model and the predictor model. We aim to evaluate the faithfulness of attribution methods for the predictor model. Due to the large output space of LMs, we use contrastive explanations proposed by Yin and Neubig (2022) that measure the attribution of input tokens for a contrastive model decision. Contrastive attributions try to identify the most important tokens that led the model to predict the target $y_{t}$ instead of a foil $y_{f}$. Then we use a separate editor model to change these important tokens to generate counterfactuals, i.e. generating examples that will make the original predictor model more likely to predict the foil.
The protocol to evaluate attributions consists of two phases. The first phase involves creating the editor that can generate counterfactuals. In the second phase of evaluating attributions, we use the editor and predictor together to determine what percentage of tokens the editor needs to change in order to flip the predictor model’s prediction. Figure 2 demonstrates the seconds phase.
Figure 1: Prompting techniques used for counterfactual generation in the second phase.
Figure 2: Our process of generating counterfactuals for evaluating attribution methods. The predictor (an LM), generates a label for the given text, and an attribution method specifics the most important tokens. We mask the top nof them and ask an editor (another LM) to change the label of the input text by filling in the masked tokens. If the attribution method is more faithful, then the needed nshould be a lower number.
Our approach for creating the editor is fine-tuning an autoregressive language model specifically for counterfactual generation. For fine-tuning, we add two tokens to the embedding space and the tokenizer, specifically ’<mask>’ and ’<counterfactual>’. For creating training examples for our counterfactual generator, inspired by Wu et al. (2021) and Donahue et al. (2020), we randomly mask between 5 and 50 percent of the tokens, then append each example’s label to it, e.g. positive or negative for SST-2 dataset, then append the <counterfactual>token, and lastly, we append the original unmasked example. The training example creation is shown in Figure 3.
In the second phase of evaluating attributions, First, we give a sentence to the predictor, then use an attribution method that was applied to the predictor to identify the most important tokens for the predictor’s decision-making process. We begin by replacing 10% of these most important tokens with ’<mask>’ and present the masked sentence and the foil label (The label with the second highest logit) to the editor to generate a counterfactual sentence (one that flips the prediction of the predictor model). If unsuccessful in flipping the prediction, we incrementally increase masking by 10% until we either flip the prediction or reach a masking threshold of 50%. This evaluation protocol is shown in Figure 2. The prompting technique used for counterfactual generation during this second phase is shown in Figure 1. The attribution technique that helps us identify the most critical tokens for creating counterfactuals and helps creating counterfactuals with the least amount of change in the original text, is the one that provides the most faithful representation of the predictor’s decision-making process.
Figure 3: Training example creation for fine-tuning the counterfactual generator, and one given sample.
4 Experimental Setup
4.1 Datasets
Three datasets are used for evaluating faithfulness. SST-2 (Socher et al., 2013) and IMDB (Maas et al., 2011) which are binary classification, and AG-News (Zhang et al., 2015) a four class classification dataset.
Faithfulness evaluation datasets should not have gold attribution labels because we do not want human intuition in faithfulness evaluation. We want to know how the model makes a prediction (Jacovi and Goldberg, 2020).
4.2 Models
4.2.1 Editor Models
For the editor model our method is very similar to Wu et al. (2021) which uses GPT-2 which is a decoder-only causal model for generating counterfactuals. We use three more modern decoder-only models. GPT-J-6B Wang and Komatsuzaki (2021) which we call gptj , and two sizes of Pythia. Pythia-1.4B (pythia1), and Pythia-2.8B (pythia2) Biderman et al. (2023). We fine-tune these models in accordance with Section 3. pythia1 model is fully fine-tuned but the other two are tuned using Low Rank Adaptation (LoRA) (Hu et al., 2022). We train these models for 8 epochs using dynamic masking Liu et al. (2019), which means masking each example differently in each epoch.
4.2.2 Predictor Models
We use Gemma-2b Team et al. (2024) as the predictor. We fine-tune the raw language model for the three tasks (gemma-ft). We use LoRA for fine-tuning. We also use an off-the-shelf instruct-tuned version (gemma-it) in zero-shot scenario. We compare these two versions (fine-tuned and non-fine-tuned) extensively for attribution evaluation.
4.3 Attribution Methods
Here we detail the six attribution methods that we use which are all widely used methods. We use all attribution methods in a contrastive way (Yin and Neubig, 2022). Contrastive attributions measure which features from the input make the foil token $y_{f}$ more likely and the target token $y_{t}$ less likely. We denote contrastive, target, and foil attributions by $S^{C}$, $S^{t}$, and $S^{f}$ respectively:
$S^{C}=S^{t}-S^{f}\vspace{-1.0ex}$ (1)
We use implementation of these attribution methods provided by Yin and Neubig (2022) (for Gradient $\times$ input, gradient norm and erasure) and by Captum Miglani et al. (2023) (for KernelSHAP and Integrated Gradient).
4.3.1 Gradient Norm
We can calculate attributions based on the norm of the gradient of the model prediction, with respect to the input (Simonyan et al., 2013; Li et al., 2016). Gradient with respect to feature $x_{i}$:
$g(x_{i})=\nabla_{x_{i}}q(y_{t}|x)$
Where $q(y_{t}|x)$ is the model output for token $y_{t}$ given the input $x$. The contrastive gradient:
$g^{C}(x_{i})=\nabla_{x_{i}}\left(q(y_{t}|\bm{x})-q(y_{f}|\bm{x})\right)$
We will use both norm one and norm two:
$S_{GN1}^{C}(x_{i})=||g^{C}(x_{i})||_{L1}$
$S_{GN2}^{C}(x_{i})=||g^{C}(x_{i})||_{L2}$
4.3.2 Gradient $\times$ Input
In gradient $\times$ input (gradinp) method Shrikumar et al. (2016); Denil et al. (2014), we take the dot product of the gradient with the input token embedding $x_{i}$:
$S_{GI}(x_{i})=g(x_{i})\cdot x_{i}$
By multiplying the gradient with the input embedding, we also account for how much each token is expressed in the attribution score. The Contrastive Gradient $\times$ Input is:
$S^{C}_{GI}(x_{i})=g^{C}(x_{i})\cdot x_{i}$
4.3.3 Erasure
Erasure-based methods measure the importance of each token by erasing it and seeing the effect on the model output (Li et al., 2017). This is achieved by taking the difference of model output with the full input $x$ and part of the input zeroed out, $x_{\neg i}$:
$S_{E}^{t}(x_{i})=q(y_{t}|x)-q(y_{t}|x_{\neg i})$
For the contrastive case:
$S^{C}_{E}=(q(y_{t}|x)-q(y_{t}|x_{\neg i}))-(q(y_{f}|x)-q(y_{f}|x_{\neg i}))$
4.3.4 KernelSHAP
KernelSHAP (Lundberg and Lee, 2017) explains the prediction of a classifier $q$ by learning a linear model $\phi$ locally around each prediction. The objective function of KernelSHAP constructs an explanation that approximates the behavior of $q$ accurately in the neighborhood of $x$. More important features have higher weights in this linear model $\phi$. Let $Z$ be a set of $N$ randomly sampled perturbations around $x$:
$S_{\phi}^{t}=\underset{\phi}{\arg\min}\sum_{z\in Z}[q(y_{t}|z)-\phi^{T}z]^{2}\pi_{x}(z)\vspace{-1.0ex}$ (2)
KernelSHAP uses a kernel $\pi_{x}$ that satisfies certain principles when input features are considered agents of a cooperative game in game theory. We use equation 2 in a contrastive way. First we normalize $S_{\phi}^{t}$ and $S_{\phi}^{f}$ by dividing to their $L2$ norm and then subtracting:
$S_{\phi}^{C}=\frac{S_{\phi}^{t}}{||S_{\phi}^{t}||}-\frac{S_{\phi}^{f}}{||S_{\phi}^{f}||}\vspace{-1.0ex}$ (3)
4.3.5 Integrated Gradients
Integrated Gradients (IG) (Sundararajan et al., 2017) is a gradient-based method which addresses the problem of saturation: gradients may get close to zero for a well-fitted function. IG requires a baseline $\mathbf{b}$ as a way of contrasting the given input with information being absent. For input $i$, we compute:
$S^{t}_{IG}=\frac{1}{m}\sum_{k=1}^{m}\nabla_{x_{i}}q\Big{(}y_{t}\Big{|}b+\frac{k}{m}(x-b)\Big{)}\!\cdot\!(x_{i}\!-\!b_{i})\!$ (4)
That is, we average over $m$ gradients, with the inputs to $f_{t}$ being linearly interpolated between the baseline and the original input $x$ in $m$ steps. We then take the dot product of that averaged gradient with the input embedding $\mathbf{x}_{i}$ minus the baseline.
We use zero vector baseline Mudrakarta et al. (2018), and 5 steps. The contrastive case becomes:
$S_{IG}^{C}=\frac{S_{IG}^{t}}{||S_{IG}^{t}||}-\frac{S_{IG}^{f}}{||S_{IG}^{f}||}\vspace{-1.0ex}$ (5)
Editor gradnorm1 Erasure KernelSHAP
gemma-ft gemma-it gemma-ft gemma-it gemma-ft gemma-it
pythia1 (ours) 1.1 1.4 1.3 1.6 0.7 1.7
pythia2 (ours) 0.4 2.6 0.9 1.3 0.8 2.3
gptj (ours) 0.7 8.3 2.0 10.9 0.9 6.4
erase 0.3 19.9 2.3 32.8 0.6 81.4
unk 0.6 97.5 1.8 97.3 1.3 99.8
mask 0.0 94.8 0.5 93.3 0.0 98.5
att-zero 0.1 80.9 0.1 62.6 0 74.1
Table 1: OOD percentage when our counterfactual editor models generate samples, compared to other replacement methods (erase, unk, mask, and att-zero methods). This represents the percentage of corrupted examples that fall outside the 99th percentile of the NLL of the original sentences in the SST-2 dataset (lower is better). Scenarios with very high OOD percentages are highlighted.
Attribution method SST-2 IMDB AG-News
pythia1 pythia2 gptj pythia1 pythia2 gptj pythia1 pythia2 gptj
gradnorm1 33.5 34.8 32.2 29.1 30.3 32.4 42.9 45.1 44.0
gradnorm2 33.4 35.6 32.6 31.0 30.5 32.4 42.6 44.4 43.9
gradinp 40.5 41.8 40.8 36.1 36.3 36.5 43.1 44.6 42.2
erasure 35.5 36.6 33.4 32.7 32.7 34.4 42.0 42.7 43.0
IG 45.7 45.8 43.7 43.3 44.3 42.5 43.8 46.7 44.0
KernelSHAP 44.1 45.9 44.9 44.0 43.3 44.2 44.0 46.5 44.3
Random 44.6 46.0 44.3 43.8 42.7 43.2 44.0 46.0 44.0
Table 2: The mean percentage needed to mask to achieve flipping Gemma-ft’s label or reaching 50 percent masking in 200 examples of evaluation split in SST-2, IMDB, and AG-News datasets (lower is better). pythia1, pythia2, and gptj models are used to fill the masks and generate counterfactuals.
Attribution method SST-2 IMDB AG-News
pythia1 pythia2 gptj pythia1 pythia2 gptj pythia1 pythia2 gptj
gradnorm1 41.4 42.0 40.3 42.1 42.3 44.0 46.6 46.8 40.0
gradnorm2 41.5 42.2 40.6 42.3 42.6 43.6 46.6 46.9 39.9
gradinp 42.9 43.4 43.2 41.2 41.4 42.2 45.8 45.3 39.2
erasure 40.8 41.5 41.0 43.1 42.8 43.8 45.0 45.5 39.5
IG 44.7 44.0 45.4 43.4 43.2 44.0 45.7 45.3 38.6
KernelSHAP 43.6 43.5 44.0 43.7 42.9 43.9 46.1 45.3 37.2
Random 44.8 44.7 44.3 45.4 46.1 45.8 44.9 45.2 39.2
Table 3: The mean percentage needed to mask to achieve flipping Gemma-it’s label or reaching 50 percent masking in 200 examples of evaluation split in SST-2, IMDB, and AG-News datasets (lower is better). pythia1, pythia2, and gptj models are used to fill the masks and generate counterfactuals.
Figure 4: In the top matrix, the average correlation of attribution ranks for the fine-tuned predictor is presented. The middle matrix shows the average correlation of attribution ranks when the predictor is an off-the-shelf model. The bottom matrix displays the difference between the first two, indicating that when using editors, the difference in correlation between the two types of predictors is near zero. However, using other replacement methods results in inconsistent correlations across the two scenarios, likely because these methods produce OOD text for the instruct-tuned model.
5 Results and Discussion
5.1 The Out-of-Distribution problem
Why should we use counterfactuals instead of erasing the important tokens or replacing important tokens with unimportant tokens?
First, we show that our counterfactual generators generate in-distribution text for the predictor. Second, we show that the rankings of faithfulness of attribution methods are consistent with each other when we are using a counterfactual generator for token replacement but these rankings differ when other replacement methods are used.
To achieve our first goal, which is to show that the generated counterfactuals are in-distribution, we use an OOD detection technique to measure what percentage of our generated inputs are OOD. Prominent methods to detect OOD use a threshold and consider anything higher than that threshold as OOD (Chen et al., 2023). To calculate the threshold, for each dataset, we measure the negative log-likelihood (NLL) of the 200 original examples using different predictors—fine-tuned and instruct-tuned—and consider the 99th percentile of the NLLs as the threshold for being OOD. The reason we use NLL to detect OOD is that the type of shift we want to detect is background shift. We can classify the types of OOD data as either semantic or background shift (Arora et al., 2021). Semantic features have a strong correlation with the label, and semantic shift happens when we encounter unseen classes at test time, while background features consist of population-level statistics that do not depend on the label and focus on the style of the text.
In the evaluation of faithfulness by corrupting the input, we are not introducing new labels or classes. We are changing the style of text; therefore, we want to detect background shift. There are two common types of OOD detection methods: calibration and density estimation. Density estimation methods, e.g., perplexity (PPL), outperform calibration methods under background shifts, while the opposite is true under semantic shift. We will use NLL, which is a very similar notion to PPL. An attribution method shows us which tokens are important, and we will replace those tokens in four ways: (i) using an editor to replace the token (our method), (ii) using a token that is considered unimportant semantically (<unk> token and <mask> token), (iii) erasing the tokens, and (iv) zeroing out the attention mask for important tokens without altering the text itself (att-zero). The baselines (ii) through (iv) are similar to previous work (Hase et al., 2021). In Table 1, it is shown that for both fine-tuned and instruct-tuned predictors, the generated counterfactuals are mostly in-distribution. In Table 1, we show the results for the SST-2 dataset and three attribution methods; other attribution methods and datasets are shown in Appendix A. Each number in the table represents the average of five amounts of replacement (10% to 50%) and 200 examples from evaluation sets.
Chen et al. (2023) demonstrate that fine-tuning renders a model insensitive to non-semantic shift. Their research indicates that fine-tuning eliminates pre-trained task-agnostic knowledge about general linguistic properties, which are crucial for detecting non-semantic shift. Our findings align with these observations. When a predictor is fine-tuned for a specific classification task, such as sentiment analysis on the SST-2 dataset, it is optimized to assign the highest probability to the positive token for examples with positive sentiment. Consequently, this fine-tuned model becomes less sensitive to input corruptions. In our experiment, regardless of the replacement method employed, the resulting input tends to remain in-distribution for the fine-tuned predictors. As evidenced in Table 1, under the Gemma-ft columns, the percentage of out-of-distribution (OOD) examples approaches zero. In contrast, Gemma-it, an off-the-shelf model which is not optimized for a specific dataset, exhibits different behavior. When subjected to various input modifications—such as replacing important tokens with semantically neutral ones (e.g., <unk> or <mask> tokens), completely removing tokens, or zeroing out the attention mask for important tokens without altering the text itself—the Gemma-it predictor frequently categorizes these modified inputs as OOD. This disparity in behavior between fine-tuned and off-the-shelf models underscores the impact of task-specific optimization on a model’s sensitivity to input perturbations. Notably, the examples remain in-distribution when modified by the counterfactual generator, even for the instruct-tuned predictor. This observation demonstrates that when we don’t want to change the predictor model and prefer to use an off-the-shelf model as our predictor, using a counterfactual generator is helpful in evaluating the faithfulness of attribution methods.
To achieve our second goal, which is to demonstrate the consistency of the faithfulness rankings of attribution methods regardless of the counterfactual generator used, and the lack of consistency when another replacement method is applied, we use Spearman’s rank correlation, which has been used in previous works (Rong et al., 2022).
We rank the attribution methods’ mask percentage needed to flip the label for each example for all seven replacement methods (our three editors, Erase, <unk>, <mask>, and att-zero), get the correlation of these ranks with each other, and then take the average over 200 examples. We show this for the SST-2 dataset in Figure 4. Other datasets yield similar results and are shown in Appendix B. In the top correlation matrix of Figure 4, these average correlations are shown for the fine-tuned predictor. For the fine-tuned predictor, all replacement methods have high average correlation with each other. The middle matrix in Figure 4 shows these correlations when the predictor models are off-the-shelf instruct-tuned models. For the off-the-shelf predictor, only when a counterfactual generator is used do the ranks have high correlation with each other; other replacement methods have low correlation with counterfactual generators. This is likely because when using an instruct-tuned predictor, replacement methods other than counterfactual generators create OOD inputs. The figure shows the rank of explanations using editors has a higher correlation with unk/erase when the predictor model is fine-tuned. The bottom matrix of 4 is the difference between the first and second matrices. It indicates that the correlation difference between fine-tuned and instruct-tuned predictors is near zero when using editors as the replacement method, but the difference with other corruption methods (unk/erase/mask/att-zero) is significant. This suggests that when evaluating explanations on an off-the-shelf instruct-tuned model, it is crucial not to use corrupted OOD text.
5.2 Analysis of FI methods
In Tables 2 and 3, we show the average masking percent needed (the average percentage of tokens the counterfactual generator should change) to flip the label for fine-tuned and instruct-tuned predictor models, respectively. The masking percent is highly correlated with flip-rate—what percentage of labels each counterfactual generator is able to flip by changing the corrupted tokens. In Appendix C, we show flip-rate for both fine-tuned and instruct-tuned predictor models. Attribution methods that are able to flip the labels with less mask percent (i.e. less change) are also able to flip more labels.
For the fine-tuned predictor (Table 2), gradient norm methods consistently outperform others on the SST-2 and IMDB datasets. In contrast, for AG-News, the Erasure method consistently performs the best or near the best. Our results indicate that straightforward methods, such as gradnorm1, gradnorm2, and Erasure, consistently deliver superior performance regardless of the editor used.
For the instruct-tuned predictor (Table 3), the Erasure method yields the best results for the SST-2 dataset, while gradinp demonstrates the best performance on the IMDB dataset. However, no attribution method consistently outperforms random selection for the AG-News dataset. Overall, these findings suggest that attribution methods are less effective when the model is not fine-tuned for the specific task, indicating the need for cautious application of these methods to pretrained and instruct-tuned language models.
6 Conclusion
In this work, we designed a faithfulness evaluation protocol based on counterfactual generation. We demonstrated that the efficacy of attribution methods varies between models fine-tuned on our specific dataset and off-the-shelf instruct-tuned models. We showed that counterfactual generators are effective for evaluating feature attribution because they can produce mostly in-distribution text for the predictor model. This approach allows us to separate the evaluation of the model from the evaluation of the attribution method, as the examples used are mostly in-distribution. We also found high consistency between different counterfactual generators and a lack of consistency with other replacement methods, highlighting the importance of being in-distribution, particularly when evaluating attributions on off-the-shelf models. Finally, we used our protocol to compare different attribution methods.
7 Limitations
Our work is limited in several aspects: First, we rely on generating counterfactuals, for which a strong generative model is needed. Generating counterfactuals, especially for long sequences, is computationally expensive. Second, the counterfactual generator may unintentionally know the artifacts and shortcuts used by the predictor to flip the label, and this could limit the intended application of our approach. Third, we evaluated our protocol for classification datasets; evaluating it for other tasks like translation is left for future work.
Appendix A
Table 4 is the OOD percentages for other attribution methods in SST-2 dataset that were not in Table 1. Tables 5 and 6 show OOD percentages in AG-News dataset.
Editor gradnorm2 gradinp integrated gradient
gemma-ft gemma-it gemma-ft gemma-it gemma-ft gemma-it
pythia1 (ours) 1.4 1.2 1.3 1.5 1.1 0.7
pythia2 (ours) 0.4 2.6 0.8 1.6 1.0 0.8
gptj (ours) 0.8 8.2 2.4 5.2 0.4 8.7
erase 0.5 21.9 1.1 83.8 1.4 77.4
unk 0.6 98.4 1.6 100.0 3.4 99.3
mask 0.0 95.8 0.2 99.5 0.4 97.8
att-zero 0.1 80.8 0.0 70.6 0 65.8
Table 4: OOD percentage when our counterfactual editor models generate samples, compared to other replacement methods (erase, unk, mask, and att-zero methods). This is the percentage of corrupted examples that are out of the 99th percentile of the NLL of the original sentences in SST-2 dataset (lower is better). The scenarios with very high numbers of OODs are highlighted.
Editor gradnorm1 Erasure KernelSHAP
gemma-ft gemma-it gemma-ft gemma-it gemma-ft gemma-it
pythia1 (ours) 2.4 5.0 4.1 4.9 1.6 4.5
pythia2 (ours) 2.4 5.2 3.1 4.9 2.0 6.9
gptj (ours) 1.0 5.0 1.5 4.9 0.9 7.8
erase 3.5 11.5 8.8 46.3 2.3 74.9
unk 1.0 98.2 3.8 99.7 1.3 99.9
mask 0.7 85.7 5.2 96.7 0.5 98.2
att-zero 4.9 79.1 3.2 57.4 0.8 62.6
Table 5: OOD percentage when our counterfactual editor models generate samples, compared to other replacement methods (erase, unk, mask, and att-zero methods). This is the percentage of corrupted examples that are out of the 99th percentile of the NLL of the original sentences in AG-News dataset (lower is better). The scenarios with very high numbers of OODs are highlighted.
Editor gradnorm2 gradinp integrated gradient
gemma-ft gemma-it gemma-ft gemma-it gemma-ft gemma-it
pythia1 (ours) 2.2 5.0 1.4 4.5 1.6 4.5
pythia2 (ours) 2.4 5.2 2.3 5.5 1.3 5.9
gptj (ours) 1.2 5.0 2.0 5.1 1.2 6.4
erase 3.3 11.6 2.6 60.3 1.9 55.5
unk 0.9 98.0 1.7 99.3 1.3 99.9
mask 0.8 86.4 1.5 94.1 0.5 98.2
att-zero 5.1 78.9 1.9 58.7 0.6 50.8
Table 6: OOD percentage when our counterfactual editor models generate samples, compared to other replacement methods (erase, unk, mask, and att-zero methods). This is the percentage of corrupted examples that are out of the 99th percentile of the NLL of the original sentences in AG-News dataset (lower is better). The scenarios with very high numbers of OODs are highlighted.
Appendix B
Figures 5 and 6 show the difference of correlations in IMDB and AG-News datasets respectively.
Figure 5: The difference
Figure 6: The difference
Appendix C
Tables 7 and 8 show flip-rates for fine-tuned and instruct-tuned datasets respectively.
Attribution method SST-2 IMDB AG-News
pythia1 pythia2 gptj pythia1 pythia2 gptj pythia1 pythia2 gptj
gradnorm1 67.5 63.5 72.5 78.5 76.5 68.8 22.0 18.5 17.0
gradnorm2 66.5 61.0 71.0 75.5 80.5 67.5 22.0 19.5 17.0
gradinp 32.0 31.0 33.0 54.5 50.5 49.0 21.5 21.0 23.0
erasure 56.5 47.5 56.5 59.0 61.5 55.5 24.5 24.0 23.0
IG 18.5 17.5 23.5 33.5 31.5 41.0 16.5 14.5 21.0
KernelSHAP 22.0 17.5 22.5 30.5 34.0 31.0 18.0 13.5 18.5
Random 22.5 18.5 23.5 33.5 38.5 35.0 18.5 15.5 18.5
Table 7: The mean percentage of success in flipping Gemma-ft’s label in 200 examples of evaluation split in SST-2, IMDB, and AG-News datasets (higher is better).
Attribution method SST-2 IMDB AG-News
pythia1 pythia2 gptj pythia1 pythia2 gptj pythia1 pythia2 gptj
gradnorm1 34.5 31.0 33.5 37.5 36.0 44.0 12.0 10.0 18.0
gradnorm2 34.0 31.5 33.5 37.0 35.0 43.5 12.0 10.0 18.5
gradinp 27.0 25.0 28.0 38.5 36.5 62.0 18.5 18.5 17.0
erasure 29.5 26.5 32.5 29.5 26.5 46.0 17.0 16.0 21.5
IG 23.5 24.5 23.0 37.0 38.0 61.0 17.5 18.5 12.0
KernelSHAP 28.0 26.0 21.5 34.0 36.0 59.5 17.5 18.0 19.0
Random 21.0 23.0 18.5 30.0 30.5 55.5 20.5 19.0 17.0
Table 8: The mean percentage of success in flipping Gemma-it’s label in 200 examples of evaluation split in SST-2, IMDB, and AG-News datasets (higher is better).